OpenStudy (anonymous):

Medal+Fan Sal’s and Carrie’s Ages Sal 3 5 8 13 15 Carrie 7 9 12 17 19 (1 point) s = c + 4 s = 4c c = s + 4 c = 4s please explain to me :(

OpenStudy (texaschic101):

c = s + 4 when Sal is 3, Carrie is 7 c = s + 4 7 = 3 + 4 7 = 7 (correct) when Sal is 5, Carrie is 9 c = s + 4 9 = 5 + 4 9 - 9 (correct) and when you check the other numbers, you will see that they all follow the same patter of : c = s + 4 understand ?

OpenStudy (anonymous):

yes I see